Interestingly, when the value of G was first measured, people all over the world were very excited because the mass of the earth can now be measured through Newton's law of universal gravitation. Using the radius of the earth, r = 6.40 X10 m, calculate its mass Consider a 1 Kg ball on the floor.

1
Expert's answer
2022-01-10T14:20:15-0500

Weight of an object on Earth =

Gravitational force exerted on an object by the Earth

"mg=\frac{GMm}{R^2}"

"g=\frac{GM}{R^2}"

"M_e=\frac{gr^2}{G}"

Put value


"M_e=\frac{9.81\times(6400000)^2}{6.67\times10^{-11}}=6.024\times10^{24}Kg"